Pleaser list all steps! Thank you so much! Using it to study.

A project requires an initial investment of $100,000 and is expected to produce a cash inflow before tax of $27,700 per year for five years. Company A has substantial accumulated tax losses and is unlikely to pay taxes in the foreseeable future. Company B pays corporate taxes at a rate of 40% and can depreciate the investment for tax purposes using the five-year MACRS tax depreciation schedule. Suppose the opportunity cost of capital is 11%. Ignore inflation.

a. Calculate project NPV for each company. (Negative answers should be indicated by a minus sign. Do not round intermediate calculations. Round your answers to the nearest whole dollar amount.)

NPV
Company A $
Company B $

b-1. What is the IRR of the after-tax cash flows for each company? (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places.)

IRR
Company A $ %
Company B $ %

b-2. What does comparison of the IRRs suggest is the effective corporate tax rate? (Do not round intermediate calculations. Enter your answer as a percent rounded to 1 decimal place.)

Effective tax rate             %
Answer #1

✓ fic =IRR(B4:39) D C E F G PVIF@11% mano B11 X A B 1 Company A: PV of aftertax 2 Year cash flows 3 A 4 0 $ - 100,000 5 1 $ 2

a) Company A Company B $2,376 $8,361 b-1) Company A Company B 11.94% 7.54% b-2) Effective tax rate 36.9% =1-7.54%/11.94%